OpenStudy (anonymous):

1/(2x-y) - 2/(x+2y).... The book answer is (4y-3x)/(2x^2+3xy-2y^2).. can someone show me how they got to that answer?

OpenStudy (anonymous):

yes use \[\frac{a}{b}-\frac{c}{d}=\frac{ad-bc}{bc}\] with lots of parentheses and lots of distributive law and lots of combine like terms

OpenStudy (anonymous):

hmmm I am understanding on how to use the formula it but i am not getting the same results? Can you show me how to do it so I can check it against my work?

OpenStudy (anonymous):

sure denominator is \[(2x-y)(x+2y)=x^2+3xy-2y^2\] by multiplying out

OpenStudy (anonymous):

numerator is \[1(x+2y)-2(2x-y)=x+2y-4x+2y=4x-3y\]

OpenStudy (anonymous):

by distributive law plus combine like terms

OpenStudy (anonymous):

Okay I just made some distribution errors...Thanks I got it now! :)